OpenStudy (anonymous):

Given parallelogram LMNO with the diagonals intersecting at point P. If MP = 2x - 1 and OP = 3x - 4, find OP

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

10

OpenStudy (anonymous):

so, MP = OP 2x-1 = 3x-4 2x-3x = -4+1 -x = -3 x=3 then, MP = 2x-1 = 2*3 -1= 6-1 = 5 MO = 2xMP= 2x5 =10 ans is D) .10

OpenStudy (anonymous):

can i get a medal plz

OpenStudy (anonymous):

@leeanne22

OpenStudy (anonymous):

thanks ! @dominique5

ganeshie8 (ganeshie8):

like the name dominique - its the main character name in my fav novel fountainhead :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

thx lol @ganeshie8

ganeshie8 (ganeshie8):

you should read fountainhead somtime... when you're free lol

OpenStudy (anonymous):

ok i will

ganeshie8 (ganeshie8):

:) let me warn you its 1000 + pages... and it you wil get addicted once u start... till you complete it !

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

The Fountainhead is a 1949 American film directed by King Vidor, based on the best-selling book of the same name by Ayn Rand, who wrote the screenplay adaptation.

ganeshie8 (ganeshie8):

:D its a very good read

ganeshie8 (ganeshie8):

dont watch the film... film was a disaster lol. novel is 1000000 times better than film ! they've fully spoled the movie... eventhough i watched the film many times, i dont recommend it... :) cya gtg